A study at Harvard found that 44% of their students binge drink. You are at Princeton doing a similar study. How many people do you need to interview to get your interval width down to 10 percentage points? For a 95% CI, what is that sample size?  
Answer #1

for an  interval width of 10 %, margin of error =10/2 =5% ~ 0.05

here margin of error E = 0.0500
for95% CI crtiical Z          = 1.96
estimated proportion=p= 0.4400
required sample size n =         p*(1-p)*(z/E)2= 379